# LLM-BT: Back-Translation as a Framework for Terminology Standardization and Dynamic Semantic Embedding

### v1

@@ Abstract
The rapid proliferation of technical terminology in English has posed major challenges to traditional expert-driven standardization mechanisms, particularly in fast-evolving fields such as artificial intelligence and quantum computing. Manual approaches struggle to meet the growing demand for multilingual consistency in scientific communication. To address this, we propose LLM-BT, a back-translation framework powered by large language models (LLMs), designed to automate terminology verification and standardization via cross-lingual semantic alignment. This work contributes three main innovations: (1) Term-Level Consistency Validation: Through the back-translation path English $\rightarrow$ intermediate language $\rightarrow$ English, LLM-BT achieves high terminology consistency across major models (e.g., GPT-4, DeepSeek, Grok). Case studies on terms like residual learning and lecanemab demonstrate over 90% exact or semantic match rates, offering both linguistic and algorithmic feasibility. (2) Multi-Path Verification Workflow: We introduce a novel “Retrieve-Generate-Verify-Optimize” pipeline that integrates both serial (e.g., EN $\rightarrow$ ZHcn $\rightarrow$ ZHtw $\rightarrow$ EN) and parallel (e.g., EN $\rightarrow$ Chinese/Portuguese $\rightarrow$ EN) BT routes. Evaluations using BLEU and term-level accuracy indicate strong cross-lingual robustness (BLEU $>$ 0.45; Portuguese term accuracy: 100%). (3) Back-Translation as Semantic Embedding: We conceptualize BT as a form of interpretable, dynamic semantic embedding. Intermediate translations reveal latent trajectories in meaning construction and alignment. Unlike static vector embeddings, LLM-BT offers transparent, path-based embeddings shaped by model evolution in compute, context and knowledge. In summary, LLM-BT redefines back-translation from a passive evaluation tool into an active engine for multilingual terminology standardization. It facilitates a human-AI collaboration paradigm: machines ensure semantic fidelity, while humans shape cultural and disciplinary interpretation. This approach offers a critical infrastructure for multilingual terminology governance across scientific, medical and technological fields, supporting global discourse in a wide range of languages and knowledge systems.
@@ 1 Introduction
National and international standardization bodies have acknowledged this urgency. For example, the China National Committee for Terminology in Science and Technology has established over 160 term systems across foundational sciences and applied domains through collaborative disciplinary committees [Zhiwei, 2004; Yang et al., 2023]. Nonetheless, current frameworks face significant limitations in the digital era: expert reviews typically require 12 to 18 months; update cycles are rigid and reactive; and mechanisms for capturing emerging terms from sources such as arXiv or ISO standards remain underdeveloped. In parallel, China’s advancements in 5G/6G, quantum technologies and AI demand efficient mechanisms for both internal localization and external dissemination of scientific terminology [Cao, 2025].
To address these challenges, this study proposes LLM-BT (Large Language Model Back-Translation), an automated terminology standardization framework that leverages multilingual back-translation powered by large language models (LLMs). The associated sub-module, LLM-BT-Terms, utilizes an English $\rightarrow$ intermediate language $\rightarrow$ English path to assess semantic consistency and translation stability, without relying on precompiled glossaries. This allows for the generation of candidate standard terms, term alignment suggestions and human-in-the-loop verification. Compared with traditional approaches, LLM-BT offers high automation, real-time adaptability and strong support for multilingual standardization tasks [Weigang and Brom, 2025].
@@
    The LLM-BT framework is proposed and evaluated in terms of terminology consistency. Experiments across simplified Chinese, traditional Chinese, Japanese and Portuguese back-translation paths show over 90% consistency, supporting its feasibility in real-world multilingual scenarios.
@@
By reframing terminology as a dynamic and verifiable construct, this work transitions term handling from passive translation to proactive standardization. LLM-BT provides a blueprint for human-AI collaboration: LLMs ensure semantic fidelity, while human experts guide contextual adaptation and cultural nuance. As such, this framework contributes to the long-term vision of establishing robust multilingual terminology infrastructure in the era of generative AI.
@@ 5.4 LLM-BT as Interpretable Semantic Embedding
Modality support Text-only Extensible to image/audio paths (e.g., EN $\rightarrow$ IMG $\rightarrow$ EN)
@@ 6.3 Semantic Embedding Differences in Simplified and Traditional Chinese Processing
In our multilingual back-translation experiments, a notable trend emerged: Traditional Chinese (ZHtw) outperforms Simplified Chinese (ZHcn) in term consistency, information retention and semantic matching. Across metrics such as Exact Match Rate (EMR), Semantic Match Rate (SMR), Information Retention Score (IRS) and Term Divergence Index (TDI), ZHtw consistently surpasses ZHcn, demonstrating greater term stability.
@@
Third, experimental data show that the EN-ZHtw-EN back-translation path outperforms EN-ZHcn-EN across term consistency metrics, closely resembling the performance of the Japanese (JA) path. This supports our hypothesis that ZHcn term recommendations should reference ZHtw practices.
These findings validate the sensitivity of the LLM-BT-Term method to intermediate language selection and highlight optimization potential for ZHcn term translation, particularly in term consistency and structural fidelity. Future research should focus on enhancing ZHcn term standardization, developing high-quality term knowledge bases, creating term-augmented models and improving the professionalism and stylistic consistency of ZHcn corpora in LLM training.
@@ 7 Conclusion
7.1 Challenges and Strategies for Back-Translation Methods

### v2

# LLM-BT-Terms: Back-Translation as a Framework for Terminology Standardization and Dynamic Semantic Embedding

@@ Abstract
The rapid growth of English technical terms challenges traditional expert-driven standardization, especially in fast-evolving fields like AI and quantum computing. Manual methods struggle to ensure multilingual consistency. We propose LLM-BT, a back-translation framework powered by large language models (LLMs) to automate terminology verification and standardization via cross-lingual semantic alignment. Our contributions are: (1) Term-Level Consistency Validation: Using English $\rightarrow$ intermediate language $\rightarrow$ English back-translation, LLM-BT achieves high term consistency across models (e.g., GPT-4, DeepSeek, Grok), with case studies showing over 90% exact or semantic matches. (2) Multi-Path Verification Workflow: A novel “Retrieve–Generate–Verify–Optimize” pipeline integrates serial (e.g., EN $\rightarrow$ ZHcn $\rightarrow$ ZHtw $\rightarrow$ EN) and parallel (e.g., EN $\rightarrow$ Chinese/Portuguese $\rightarrow$ EN) BT routes. BLEU and term accuracy indicate strong cross-lingual robustness (BLEU $>$ 0.45; Portuguese accuracy 100%). (3) Back-Translation as Semantic Embedding: BT is conceptualized as dynamic semantic embedding, revealing latent meaning trajectories. Unlike static embeddings, LLM-BT provides transparent path-based embeddings shaped by model evolution. LLM-BT transforms back-translation into an active engine for multilingual terminology standardization, enabling human–AI collaboration: machines ensure semantic fidelity, humans guide cultural interpretation. This infrastructure supports terminology governance across scientific and technological fields worldwide.
@@ 1 Introduction
National and international standardization bodies have acknowledged this urgency. For example, the China National Committee for Terminology in Science and Technology has established over 160 term systems across foundational sciences and applied domains through collaborative disciplinary committees [Feng, 1997; Liu, 2015; Yang et al., 2023]. Nonetheless, current frameworks face significant limitations in the digital era: expert reviews typically require 12 to 18 months; update cycles are rigid and reactive; and mechanisms for capturing emerging terms from sources such as arXiv or ISO standards remain underdeveloped. In parallel, China’s advancements in 5G/6G, quantum technologies and AI demand efficient mechanisms for both internal localization and external dissemination of scientific terminology [Cao, 2025].
In Brazil, the development of archival terminology dictionaries involves collaboration among archivists, librarians, linguists, and public institutions to create a standardized technical vocabulary that reflects Brazilian Portuguese while aligning with international standards, such as those of the International Council on Archives (ICA). The Brazilian Standards Association (ABNT) and the Brazilian Archivists Association (AAB) lead terminology standardization efforts, while the National Archives and the São Paulo State Public Archives contributed to regional and national dictionaries in the 1980s and 1990s, adapting English terms to local practices [Camargo and Bellotto, 2005]. In 1997, the Ibero-American Group for Administrative Archives Processing fostered term convergence among Brazil, Portugal, and Spanish-speaking countries, influencing English translations. Terms without direct equivalents, such as “provenance” (“proveniência” or “respect for fonds”), require contextual decisions, documented in dictionaries.
To address these challenges, this study proposes LLM-BT (Large Language Model Back-Translation), an automated terminology standardization framework that leverages multilingual back-translation powered by large language models (LLMs). The associated sub-module, LLM-BT-Terms, utilizes an English $\rightarrow$ intermediate language $\rightarrow$ English path to assess semantic consistency and translation stability, without relying on precompiled glossaries, see 1. This allows for the generation of candidate standard terms, term alignment suggestions and human-in-the-loop verification. Compared with traditional approaches, LLM-BT offers high automation, real-time adaptability and strong support for multilingual standardization tasks [Weigang and Brom, 2025]. The LLM-BT-Terms model can also support translation standardization by using the EN→PTbr→PTpt→EN scheme to validate semantic and cultural consistency.
@@
    The LLM-BT-Terms framework is proposed and evaluated in terms of terminology consistency. Experiments across simplified Chinese, traditional Chinese, Japanese and Portuguese back-translation paths show over 90% consistency, supporting its feasibility in real-world multilingual scenarios.
@@
By reframing terminology as a dynamic and verifiable construct, this work transitions term handling from passive translation to proactive standardization and leverages the LLM-BT-Terms framework to address such challenges, offering a scalable solution for cross-lingual terminology alignment, as demonstrated in the Brazilian and Chinese context. LLM-BT-Terms provides a blueprint for human-AI collaboration: LLMs ensure semantic fidelity, while human experts guide contextual adaptation and cultural nuance. As such, this framework contributes to the long-term vision of establishing robust multilingual terminology infrastructure in the era of generative AI.
@@ 5.4 LLM-BT as Interpretable Semantic Embedding
Modality support Multimodal with fusion Extensible to image/audio paths (e.g., EN $\rightarrow$ IMG $\rightarrow$ EN) and other modals
@@ 6.3 Semantic Embedding Differences in Simplified and Traditional Chinese Processing
In our multilingual back-translation experiments using LLM-BT-Terms, a notable trend emerged: Traditional Chinese (ZHtw) outperforms Simplified Chinese (ZHcn) in term consistency, information retention and semantic matching. Across metrics such as Exact Match Rate (EMR), Semantic Match Rate (SMR), Information Retention Score (IRS) and Term Divergence Index (TDI), ZHtw consistently surpasses ZHcn, demonstrating greater term stability.
@@
Third, experimental data show that the EN-ZHtw-EN back-translation path outperforms EN-ZHcn-EN across term consistency metrics, closely resembling the performance of the Japanese (JA) path. This factor further supports the central hypothesis of this study: the standardization of Chinese terminology should consider the similarities and differences across Simplified Chinese (ZHcn), Traditional Chinese (ZHtw), and even Japanese corpora, aiming to unify term expressions and improve the consistency and quality of translations.
These findings validate the sensitivity of the LLM-BT-Terms method to intermediate language selection and highlight optimization potential for ZHcn term translation, particularly in term consistency and structural fidelity. Crucially, the development of the LLM-BT-Terms mechanism enables this important task by significantly reducing the manual effort traditionally required for cross-variant comparison, thus making large-scale terminology alignment practically achievable. Future research should focus on enhancing ZHcn term standardization, developing high-quality term knowledge bases, creating term-augmented models and improving the professionalism and stylistic consistency of ZHcn corpora in LLM training.
@@ 7 Conclusion
7.1 Challenges and Strategies for LLM-Back-Translation Methods